buyers in the chocolate market demand 200 units of chocolate when price is zero and decrease the purchase of chocolates by 2 units for every one unit increase in price. write down the equation of the demand function in the form of Q= f(P)


1
Expert's answer
2021-05-25T11:12:28-0400

In its standard form a linear demand equation is Q = a - bP. That is, quantity demanded is a function of price. 

"P= 0 , Q = 200 .... ( Point 1)"

"P= 1 , Q = 198 ......(point 2)"

"Q = a - bP"

When we put point 1 , we get 

"a = 200"

When we put point 2 , we get 

"198 = 200 - b"

"b = 2"

Therefore the equation becomes, 

"Q = 200 - 2P"